Content and Structure – The paper

1. (20) Has a clearly stated topic or question that it will address

2. (5) Does address the stated topic or question

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas c. (10) Transitions between paragraphs are logical

3. (10) Accurately presents facts and viewpoints of sources

4. (15) Provides sufficient textual evidence to support the argument

5. (10) Is concise, clear, and not too repetitive

6. (10) Has a clearly identifiable conclusion

Style and English Usage – The paper

1. (20) Is clear and direct in its use of language

2. (10) Is free of misspellings

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns)

4. (10) Lacks incomplete sentences

5. (10) Lacks run-on or overly complicated sentences

5. (10) Uses correct punctuation

6. (15) Is free of jargon, cliches, and misused words

7. (10) Cites quotations and/or references correctly